OpenStudy (anonymous):

Ican Tackel just signed a $11.5 million, 4-year contract with an NFL team. He received a signing bonus of $2 million. He will then be paid the following amounts over his 4 year contract: $1.5 at the end of year 1; $3 million at the end of year 2; $3.5 million at the end of year 3; and $1.5 million at the end of year 4. What is the present value of his contract if money can earn 8% per year?

OpenStudy (anonymous):

I am trying to find the correct formula to use

OpenStudy (dumbcow):

PV = 2 + 1.5v +3v^2 + 3.5v^3 + 1.5v^4 v = 1/(1+i)